# The Solution to an Impulse Control Problem Motivated by Optimal Harvesting

### v1

@@ 1 Introduction
Stochastic impulse control problems arise in various fields. In the context of mathematical finance, economics and operations research, notable contributions include Harrison, Sellke and Tayor [17], Harrison and Taksar [18], Mundaca and Øksendal [24], Korn [19, 20], Bielecki and Pliska [9], Cadenillas [11], Bar-Ilan, Sulem and Zanello [6], Bar-Ilan, Perry and Stadje [5], Ohnishi and Tsujimura [25], Cadenillas, Sarkar and Zapatero [12], LyVath, Mnif and Pham [23], and several references therein. Also, models motivated by the optimal management of a natural resource have been studied by Alvarez [1, 2], Alvarez and Koskela [3] and Alvarez and Lempa [4]. In view of the wide range of applications, the general mathematical theory of stochastic impulse control is well-developed: for instance, see Lepeltier and Marchal [22], Perthame [27], Egami [15], Guo and Wu [16], Davis, Guo and Wu [13], Djehiche, Hamadène and Hdhiri [14], Belak, Christensen and Seifried [7], as well as the books by Bensoussan and Lions [8], Øksendal and Sulem [26], Pham [28], and several references therein.
@@
Beyond obtaining the complete solution to the impulse control problem that we consider, we construct a unique strong solution to the SDE (1) when the controlled process $\zeta$ is a $\beta$-$\gamma$ strategy (see Theorem 5 in Section 4). Despite the central role that such strategies play in stochastic impulse control, we are not aware of any such rigorous result. Furthermore, although such a result is rather intuitive, it turns out that it is not straightforward to establish. The proof that we develop can most easily be adapted to derive the existence of strong solutions to optimally controlled SDEs that arise in other stochastic impulse control problems, even in dimensions higher than one.
@@ 2 Formulation of the stochastic control problem
The assumption that the boundary points 0 and $\infty$ are inaccessible implies that they can be either natural or entrance ones. If 0 is a natural boundary point, then
@@
where $\Lambda$ is defined by (3) with $X$ in place of $X^{\zeta}$, and $T_{y}$ is the first hitting time of the set $\{y\}$, which is defined by
$T_{y}=\inf\left\{t\geq 0\mid\ X_{t}=y\right\},\quad\text{for }y>0.$ (15)
@@ 3 Results associated with a linear ODE
Unless stated otherwise, the results in this section hold true if the coefficients of (4) satisfy the usual Engelbert and Schmidt conditions, rather than the stronger Assumption 1, and the boundary points 00, $\infty$ are inaccessible. We start by recalling some standard results that we will need and can be found in, e.g., Lamberton and Zervos [21, Section 4]. Consider a Borel measurable function $F:\mbox{}]0,\infty[\rightarrow\mathbb{R}$ such that
@@
where we have also used (31). Using the first of these two observations, the definition (23) of $R_{\Theta}$, (37) and integration by parts, we can see that, if $\lim_{x\downarrow 0}\Theta(x)/r(x)=-\infty$, then
@@
$\lim_{x\downarrow 0}\frac{R^{\prime}_{\Theta}(x)}{\psi^{\prime}(x)}=\lim_{x\downarrow 0}\frac{\frac{\mathrm{d}}{\mathrm{d}x}\frac{R^{\prime}_{\Theta}(x)}{\varphi^{\prime}(x)}}{\frac{\mathrm{d}}{\mathrm{d}x}\frac{\psi^{\prime}(x)}{\varphi^{\prime}(x)}}=\infty,$
and the implications in (45) follow from the definition (43) of $\overline{x}$. $\Box$
@@ 4 The “$\boldsymbol{\beta}$-$\boldsymbol{\gamma}$” strategy
The process $\widetilde{W}^{\ell+1}$ defined by $\widetilde{W}_{t}^{\ell+1}=\bigl{(}W_{\tau_{\ell}+t}-W_{\tau_{\ell}}\bigr{)}{\bf 1}_{\{\tau_{\ell}<\infty\}}$ is a standard $(\mathcal{F}_{\tau_{\ell}+t})$-Brownian motion that is independent of $\mathcal{F}_{\tau_{\ell}}$ under the conditional probability measure $\operatorname{\mathbb{P}}(\cdot\mid\tau_{\ell}<\infty)$ (see Revuz and Yor [29, Exercise IV.3.21]). We denote by $\widetilde{X}^{\ell+1}$ the unique solution to the uncontrolled SDE (4) with $\widetilde{X}_{0}^{\ell+1}=\gamma$ that is driven by the Brownian motion $\widetilde{W}^{\ell+1}$ and is defined on the probability space $\bigl{(}\Omega,\mathcal{F},(\mathcal{F}_{\tau_{\ell}+t}),\operatorname{\mathbb{P}}(\cdot\mid\tau_{\ell}<\infty)\bigr{)}$. Since $(t-\tau_{\ell})^{+}$ is an $(\mathcal{F}_{\tau_{\ell}+t})$-stopping time for all $t\geq 0$,
@@
where we have time changed the Lebesgue as well as the Itô integral (see Revuz and Yor [29, Propositions V.1.4, V.1.5]). It follows that, if we define
@@ 5 The solution to the control problem
Proof. The boundedness from below of $w$ follows immediately from the conditions in (i) and (ii) of Assumption 4 and Lemma 1.(i).
@@
Proof. The equivalence (89) follows immediately from the statement related to (75) in Lemma 6 and part (II) of Lemma 8. On the other hand, the boundedness from below of $w$ follows immediately from the conditions in (i) and (ii) of Assumption 4 and Lemma 1.(i).
@@
We next consider any sequence $(\tau_{n})$ of bounded localising times for the local martingale $M^{\zeta}$. Recalling that $h$ and $w$ are both bounded from below, we use Fatou’s lemma, the monotone convergence theorem and the admissibility condition (7) to observe that (93) implies that
@@
In view of Lemma 1.(ii), if 0 is a natural boundary point, then, in all of the cases appearing in Lemmas 7-9,
@@ 6 Examples
for some constants $b$ and $\sigma>0$. Furthermore, if $r>b$ and $h$ is any strictly concave function such that
$\lim_{x\downarrow 0}h^{\prime}(x)>k(r-b)\quad\text{and}\quad\lim_{x\uparrow\infty}h^{\prime}(x)<k(r-b),$
@@
for some strictly positive constants $κ\upkappa$, $\upgamma$, $\upsigma$ and $\ell\in[1,\frac{3}{2}]$. Note that the celebrated stochastic Verhulst-Pearl logistic model of population growth arises in the special case $\ell=1$. Assumptions 1-3 hold true if $\ell\in\mbox{}]1,\frac{3}{2}]$ or if $\ell=1$ and $k\upgamma-\frac{1}{2}\sigma^{2}>0$. Furthermore, if $h$ is any concave function such that
@@
Furthermore, if $h$ is any concave function, then $\Theta$ is strictly concave and satisfies the requirements of Assumption 4. $\Box$
@@
$\lim_{x\downarrow 0}h^{\prime}(x)>k(r+\kappa)\quad\text{and}\quad\lim_{x\uparrow\infty}h^{\prime}(x)<k(r+\kappa),$
@@
for some constants $\alpha\in\mbox{]}1,2]$ and $a\in\mbox{}\bigl{]}0,\frac{3}{2}(2b+\sigma^{2})(\alpha-1)\bigr{[}$. In this case, all of the conditions in Assumption 4 hold true,
@@
$\text{and}\quad R_{\Theta}(x)=\begin{cases}x^{2}+\frac{2a}{\sigma^{2}(n-\alpha)(\alpha-m)}x^{\alpha}-\bigl{(}3+\frac{2a}{\sigma^{2}(n-m)n(n-\alpha)}\bigr{)}x,&\text{if }x\in\mbox{}]0,1[,\\ \frac{a}{r}-3+x^{-1}+\frac{2a}{\sigma^{2}(n-m)m(\alpha-m)}x^{m},&\text{if }x\geq 1.\end{cases}$
@@
$r(x)=\alpha,\quad h(x)=\begin{cases}e^{x}-1,&\text{if }x\in\mbox{}]0,1[,\\ e+e^{\gamma}-1-e^{\gamma x},&\text{if }x\geq 1,\end{cases}\quad\text{and}\quad k(x)=\kappa,\quad\text{for }x>0,$
for some constants $\gamma\in\mbox{}]0,1[$ and $\kappa\in\mbox{}\bigl{]}0,\frac{1}{2\alpha}\bigr{[}$. In this case,
@@
$\Theta(x)=\begin{cases}2\alpha\kappa-1-2\alpha\kappa x+e^{x},&\text{if }x\in\mbox{}]0,1[,\\ 2\alpha\kappa+e+e^{\gamma}-1-2\alpha\kappa x-e^{-\gamma x},&\text{if }x\geq 1,\end{cases}$
@@
$\lim_{x\downarrow 0}R_{h}(x)=\frac{1}{\alpha}\biggl{(}1-\frac{\gamma}{1-\gamma}e^{-(1-\gamma)}\biggr{)}=:\frac{1}{\alpha}f(\gamma)\quad\text{and}\quad\lim_{x\downarrow 0}\frac{R^{\prime}_{\Theta}(x)}{\psi^{\prime}(x)}=\frac{1}{\alpha}f(\gamma)-2\kappa.$
The function $f$ is strictly decreasing, $f(0)=1$ and $\lim_{\gamma\uparrow 1}f(\gamma)=-\infty$. Therefore, there exist constants $10<\gamma_{1}<\gamma_{2}<$ 1 such that $f(\gamma)\in\mbox{}]0,2\alpha\kappa[$ for all $\gamma\in\mbox{}]\gamma_{1},\gamma_{2}[$ and $f(\gamma)<0$ for all $\gamma\in\mbox{}]\gamma_{2},1[$. In view of these observations, we can see that
$\lim_{x\downarrow 0}R_{h}(x)\neq 0=\lim_{x\downarrow 0}\frac{h(x)}{r(x)}\quad\text{ for all }\gamma\in\mbox{}]0,1[\mbox{}\setminus\{\gamma_{2}\},$ (98)
which shows that (27) in Lemma 1 is not in general true if 0 is an entrance boundary point. On the other hand, Lemma 4 implies that, if $\gamma\in\mbox{}]\gamma_{1},1[$, then $0<\underline{x}<\overline{x}<\infty$ and we are in the context of Lemma 8 with $c^{\circ}=\infty$. In this context, (88) yields the expression
@@
(a) If $\gamma\in\mbox{}]0,\gamma_{2}[$, then $w(0)>0=h(0)/r(0)$ and a $\beta$-0 strategy would be strictly sub-optimal.
(b) If $\gamma\in\mbox{}]\gamma_{2},1[$, then $w(0)<0=h(0)/r(0)$ for all $c$ sufficiently large, in which case, a $\beta$-0 strategy would be optimal. $\Box$

### v2

@@ 1 Introduction
Stochastic impulse control problems arise in various fields. In the context of mathematical finance, economics and operations research, notable contributions include Harrison, Sellke and Tayor [20], Harrison and Taksar [21], Mundaca and Øksendal [32], Korn [24, 25], Bielecki and Pliska [9], Cadenillas [11], Bar-Ilan, Sulem and Zanello [7], Bar-Ilan, Perry and Stadje [6], Ohnishi and Tsujimura [33], Cadenillas, Sarkar and Zapatero [12], LyVath, Mnif and Pham [30], and several references therein. Also, impulse control models motivated by the optimal management of a natural resource have been studied by Alvarez [1, 2], Alvarez and Koskela [4] and Alvarez and Lempa [5]; singular control versions of such models have been studied by Lungu and Øksendal [29], Framstad [19], Song, Stockbridge and Zhu [40], Alvarez and Hening [3] and several references therein. In view of the wide range of applications, the general mathematical theory of stochastic impulse control is well-developed: apart from the contributions mentioned above, see also Richard [39], Stettner [41], Lepeltier and Marchal [28], Perthame [36], Egami [18], Davis, Guo and Wu [16], Djehiche, Hamadène and Hdhiri [17], Christensen [13], Helmes, Stockbridge and Zhu [22, 23], Menaldi and Robin [31], Palczewski and Stettner [35], Christensen and Strauch [14], as well as the books by Bensoussan and Lions [8], Davis [15], Øksendal and Sulem [34], Pham [37], and several references therein.
@@
The evolution of an impulse control problem’s state process is quite intuitive, provided that the corresponding uncontrolled dynamics are well-posed. For this reason, several references simply assume the existence of such processes. In the context of SDEs in $\mathbb{R}^{d}$, the state process of an impulse control problem can be derived by pasting together suitable strong solutions to the underlying uncontrolled SDE with random initial conditions (e.g., see Bensoussan and Lions [8, Section 6.1.1]). In the context of general Markov processes, the classical construction of an impulse control strategy is substantially more technical and may involve countable products of canonical spaces (e.g., see Stettner [41] and Lepeltier and Marchal [28]). If the uncontrolled state space process is a general Markov process with continuous sample paths, then comprehensive constructions of impulse control models have been derived by Helmes, Stockbridge and Zhu [23].
Impulse control problems with SDEs in $\mathbb{R}^{d}$ can be formulated as in (1)–(3). In itself such a formulation is straightforward. Indeed, an SDE in $\mathbb{R}^{d}$ such as (1) has a unique strong solution under suitable Lipschitz assumptions on $b$ and $\sigma$ for a wide class of controlled processes $\zeta$ (e.g., see Krylov [26, Theorem 2.5.7]). On the other hand, a rigorous construction of an optimally controlled process $\zeta$, such as a $\beta$-$\gamma$ strategy, is rather non-trivial. In the context of this paper, we construct a unique strong solution to the SDE (1) when the controlled process $\zeta$ is a $\beta$-$\gamma$ strategy (see Theorem 5 in Section 4). Despite the central role that such strategies play in stochastic impulse control, we are not aware of any such rigorous SDE result. Furthermore, this construction allows for a probabilistic derivation of the optimal expected discounted running reward as well as the optimal expected discounted reward from control exdenditure functionals (see (49) and (50) in Theorem 5). The construction that we make can most easily be adapted to derive the existence of strong solutions to optimally controlled SDEs that arise in other stochastic impulse control problems, even in dimensions higher than one.
@@ 2 Formulation of the stochastic control problem
The state space of the linear diffusion associated with the SDE (4) is the interval $\mathcal{I}=\mbox{}]0,\infty[$. Recall that the boundary point $p\in\{0,\infty\}$ of $\mathcal{I}$ is called inaccessible if $\operatorname{\mathbb{P}}_{x}(T_{p}<\infty)=0$ for all $x\in\mathcal{I}$ and accessible otherwise. Furthermore, if the boundary $p$ is inaccessible, then it is natural if
$\lim_{x\in\mathcal{I},\,x\rightarrow p}\operatorname{\mathbb{P}}_{x}(T_{y}<t)=0\quad\text{for all }y\in\mathcal{I}\text{ and }t>0$
and entrance otherwise, namely, if
$\lim_{x\in\mathcal{I},\,x\rightarrow p}\operatorname{\mathbb{P}}_{x}(T_{y}<t)>0\quad\text{for some }y\in\mathcal{I}\text{ and }t>0$
(e.g., see Revuz and Yor [38, Definition VII.3.9]). In these expressions, $T_{y}$ is the first hitting time of the set $\{y\}$, which is defined by
$T_{y}=\inf\left\{t\geq 0\mid\ X_{t}=y\right\},\quad\text{for }y>0.$ (7)
In Borodin and Salminen [10, II.1.6], an inaccessible boundary point is called not-exit, while a natural (resp., entrance) boundary point is called natural (resp., entrance-not-exit). Integral conditions for the classification of a boundary point $p\in\{0,\infty\}$ of $\mathcal{I}$ in terms of the scale function $p$ and the speed measure $m$ can be found in this reference.
@@
If 0 is a natural boundary point, then
@@
where $\Lambda$ is defined by (3) with $X$ in place of $X^{\zeta}$ and $T_{y}$ is defined by (7).
@@ 3 Results associated with a linear ODE
Unless stated otherwise, the results in this section hold true if the coefficients of (4) satisfy the usual Engelbert and Schmidt conditions, rather than the stronger Assumption 1, and the boundary points 00, $\infty$ are inaccessible. We start by recalling some standard results that we will need and can be found in, e.g., Lamberton and Zervos [27, Section 4]. Consider a Borel measurable function $F:\mbox{}]0,\infty[\rightarrow\mathbb{R}$ such that
@@
where we have also used (31). Using the first of these two observations, the definition (23) of $R_{\Theta}$, (34), (37) and integration by parts, we can see that, if $\lim_{x\downarrow 0}\Theta(x)/r(x)=-\infty$, then
@@
$\lim_{x\downarrow 0}\frac{R^{\prime}_{\Theta}(x)}{\psi^{\prime}(x)}=\lim_{x\downarrow 0}\frac{\frac{\mathrm{d}}{\mathrm{d}x}\frac{R^{\prime}_{\Theta}(x)}{\varphi^{\prime}(x)}}{\frac{\mathrm{d}}{\mathrm{d}x}\frac{\psi^{\prime}(x)}{\varphi^{\prime}(x)}}=\infty.$
The implications in (45) follow from this analysis and the definition (43) of $\overline{x}$. $\Box$
@@ 4 The “$\boldsymbol{\beta}$-$\boldsymbol{\gamma}$” strategy
The process $\widetilde{W}^{\ell+1}$ defined by $\widetilde{W}_{t}^{\ell+1}=\bigl{(}W_{\tau_{\ell}+t}-W_{\tau_{\ell}}\bigr{)}{\bf 1}_{\{\tau_{\ell}<\infty\}}$ is a standard $(\mathcal{F}_{\tau_{\ell}+t})$-Brownian motion that is independent of $\mathcal{F}_{\tau_{\ell}}$ under the conditional probability measure $\operatorname{\mathbb{P}}(\cdot\mid\tau_{\ell}<\infty)$ (see Revuz and Yor [38, Exercise IV.3.21]). We denote by $\widetilde{X}^{\ell+1}$ the unique solution to the uncontrolled SDE (4) with $\widetilde{X}_{0}^{\ell+1}=\gamma$ that is driven by the Brownian motion $\widetilde{W}^{\ell+1}$ and is defined on the probability space $\bigl{(}\Omega,\mathcal{F},(\mathcal{F}_{\tau_{\ell}+t}),\operatorname{\mathbb{P}}(\cdot\mid\tau_{\ell}<\infty)\bigr{)}$. Since $(t-\tau_{\ell})^{+}$ is an $(\mathcal{F}_{\tau_{\ell}+t})$-stopping time for all $t\geq 0$,
@@
where we have time changed the Lebesgue as well as the Itô integral (see Revuz and Yor [38, Propositions V.1.4, V.1.5]). It follows that, if we define
@@ 5 The solution to the control problem
Proof. The boundedness from below of $w$ follows immediately from Assumption 3, the conditions in (i) and (ii) of Assumption 4 and Lemma 1.(i).
@@
Proof. The equivalence (89) follows immediately from the statement related to (75) in Lemma 6 and part (II) of Lemma 8. On the other hand, the boundedness from below of $w$ follows immediately from Assumption 3, the conditions in (i) and (ii) of Assumption 4 and Lemma 1.(i).
@@
We next consider any sequence $(\tau_{n})$ of bounded localising times for the local martingale $M^{\zeta}$. Recalling Assumption 4.(ii) as well as the fact that $h$ and $w$ are both bounded from below, we use Fatou’s lemma, the monotone convergence theorem and the admissibility condition (8) to observe that (93) implies that
@@
In view of the second limit in (13) and Lemma 1.(ii), if 0 is a natural boundary point, then, in all of the cases appearing in Lemmas 7-9,
@@ 6 Examples
for some constants $b$ and $\sigma>0$. Furthermore, if $r>b$ and $h$ is any bounded from below strictly concave function such that
$\lim_{x\downarrow 0}h^{\prime}(x)>k(r-b)\quad\text{and}\quad\lim_{x\uparrow\infty}h^{\prime}(x)=0,$
@@
for some strictly positive constants $κ\upkappa$, $\upgamma$, $\upsigma$ and $\ell\in[1,\frac{3}{2}]$. Note that the celebrated stochastic Verhulst-Pearl logistic model of population growth arises in the special case $\ell=1$. Assumptions 1-3 hold true if $\ell\in\mbox{}]1,\frac{3}{2}]$ or if $\ell=1$ and $k\upgamma-\frac{1}{2}\sigma^{2}>0$. Furthermore, if $h$ is any bounded from below concave function such that
@@
Furthermore, if $h$ is any bounded from below concave function, then $\Theta$ is strictly concave and satisfies the requirements of Assumption 4. $\Box$
@@
$\lim_{x\downarrow 0}h^{\prime}(x)>k(r+\kappa)\quad\text{and}\quad\lim_{x\uparrow\infty}h^{\prime}(x)=0,$
@@
for some constants $\alpha\in\mbox{]}1,2]$ and $a\in\mbox{}\bigl{]}0,\frac{3}{2}(2b+\sigma^{2})(\alpha-1)\vee 3r\bigr{[}$. In this case, all of the conditions in Assumption 4 hold true,
@@
$\text{and}\quad R_{\Theta}(x)=\begin{cases}x^{2}+\frac{2a}{\sigma^{2}(n-\alpha)(\alpha-m)}x^{\alpha}-\bigl{(}3+\frac{2a\alpha}{\sigma^{2}(n-m)n(n-\alpha)}\bigr{)}x,&\text{if }x\in\mbox{}]0,1[,\\ \frac{a}{r}-3+x^{-1}+\frac{2a\alpha}{\sigma^{2}(n-m)m(\alpha-m)}x^{m},&\text{if }x\geq 1.\end{cases}$
@@
$r(x)=\alpha,\quad h(x)=\begin{cases}e^{x}-1,&\text{if }x\in\mbox{}]0,1[,\\ e-e^{\gamma+3}-1+e^{\gamma x+3},&\text{if }x\geq 1,\end{cases}\quad\text{and}\quad k(x)=\kappa,\quad\text{for }x>0,$
for some constants $\gamma<0$ and $\kappa\in\mbox{}\bigl{]}0,\frac{1}{2\alpha}\bigr{[}$. In this case,
@@
$\Theta(x)=\begin{cases}2\alpha\kappa-1-2\alpha\kappa x+e^{x},&\text{if }x\in\mbox{}]0,1[,\\ 2\alpha\kappa+e-e^{\gamma+3}-1-2\alpha\kappa x+e^{\gamma x+3},&\text{if }x\geq 1,\end{cases}$
@@
$\lim_{x\downarrow 0}R_{h}(x)=\frac{1}{\alpha}\biggl{(}1+\frac{\gamma}{1-\gamma}e^{\gamma+2}\biggr{)}=:\frac{1}{\alpha}f(\gamma)\quad\text{and}\quad\lim_{x\downarrow 0}\frac{R^{\prime}_{\Theta}(x)}{\psi^{\prime}(x)}=\frac{1}{\alpha}f(\gamma)-2\kappa.$
The function $f$ is strictly decreasing in the interval $\bigl{]}{-\infty},(1-\sqrt{5})/2\bigr{[}$, strictly increasing in the interval $\bigl{]}(1-\sqrt{5})/2,0\bigr{[}$,
$\lim_{\gamma\downarrow-\infty}f(\gamma)=1,\quad f\biggl{(}\frac{1-\sqrt{5}}{2}\biggr{)}=1-\frac{1}{2}\bigl{(}3-\sqrt{5}\bigr{)}e^{(5-\sqrt{5})/2}<0\quad\text{and}\quad f(0)=1.$
Therefore, there exist constants $(1-\sqrt{5})/2<\gamma_{1}<\gamma_{2}<0$ such that $f(\gamma)<0$ for all $\gamma\in\bigl{[}(1-\sqrt{5})/2,\gamma_{1}\bigr{[}$ and $f(\gamma)\in\mbox{}]0,2\alpha\kappa[$ for all $\gamma\in\mbox{}]\gamma_{1},\gamma_{2}[$. In view of these observations, we can see that
$\lim_{x\downarrow 0}R_{h}(x)\neq 0=\lim_{x\downarrow 0}\frac{h(x)}{r(x)}\quad\text{ for all }\gamma\in\bigl{[}(1-\sqrt{5})/2,\gamma_{1}\bigr{[}\mbox{}\setminus\{\gamma_{2}\},$ (98)
which shows that (27) in Lemma 1 is not in general true if 0 is an entrance boundary point. On the other hand, Lemma 4 implies that, if $\gamma\in\mbox{}\bigl{[}(1-\sqrt{5})/2,\gamma_{1}\bigr{[}$, then $0<\underline{x}<\overline{x}<\infty$ and we are in the context of Lemma 8 with $c^{\circ}=\infty$. In this context, (88) yields the expression
@@
(a) If $\gamma\in\mbox{}]\gamma_{1},0[$, then $w(0)>0=h(0)/r(0)$ and a $\beta$-0 strategy would be strictly sub-optimal.
(b) If $\gamma\in\mbox{}\bigl{]}(1-\sqrt{5})/2,\gamma_{1}\bigr{[}$, then $w(0)<0=h(0)/r(0)$ for all $c$ sufficiently large, in which case, a $\beta$-0 strategy would be optimal. $\Box$